OpenStudy (rizags):

Two parabolas have equations y= x^2 + ax +b and y= x^2 + cx +d, where a, b, c, and d are integers, each chosen independently by rolling a fair six-sided die. What is the probability that the parabolas will have a least one point in common?

jimthompson5910 (jim_thompson5910):

you're basically finding the probability of the two parabolas intersecting (where a does not equal c, and b does not equal d)

jimthompson5910 (jim_thompson5910):

a,,b,c,d are all whole numbers between 1 and 6 (including both 1 and 6)

OpenStudy (rizags):

how do i find the chance that a will be equal to c?

OpenStudy (rizags):

waiiit, there a 1/6 chance of a=c and 1/6 chance of b=d right?

jimthompson5910 (jim_thompson5910):

yes correct

OpenStudy (rizags):

what do i do from here though?

jimthompson5910 (jim_thompson5910):

what happens when you set the two equations y= x^2 + ax +b and y= x^2 + cx +d, equal to one another? what do you notice cancelling?

OpenStudy (rizags):

well, the x^2

jimthompson5910 (jim_thompson5910):

so we'll have ax + b = cx + d solve for x to get ??

OpenStudy (rizags):

x(a-c) = d-b

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (rizags):

well, \[\frac{ d-b }{ a-c }=x \]

jimthompson5910 (jim_thompson5910):

ok let me think for a second

jimthompson5910 (jim_thompson5910):

ok I just realized that if a = c and b = d, then we'll have 2 identical parabolas which means that every point on either parabola is on the other parabola. Since they say "AT LEAST ONE" point in common, this means that a = c and b = d is possible. For instance, if a = 2, b = 3 and a = c, b = d, then y= x^2 + ax +b and y= x^2 + cx +d turn into y= x^2 + 2x +3 and y= x^2 + 2x +3 they'll have infinitely many points in common since they are the same parabola

jimthompson5910 (jim_thompson5910):

as for generating every possible case, I'm still thinking that one over

jimthompson5910 (jim_thompson5910):

this is a tough problem, so I looked it up and found this page http://www.artofproblemsolving.com/Wiki/index.php/2012_AMC_12B_Problems/Problem_13 apparently it's part of the AMC (American Mathematics Competitions) and they're known to ask really tough questions

OpenStudy (rizags):

could you explain their solution 1 to me

jimthompson5910 (jim_thompson5910):

which part of solution 1?

jimthompson5910 (jim_thompson5910):

1296 = 6^4. There are 4 dice and 6 faces per die, so 6^4 = 1296 possible ways to roll 4 dice. they are getting 1080 from computing 5*6^3. They say "assume a is not equal to c", to that means one of the dice (either a or c) has 1 less choice to land on.

jimthompson5910 (jim_thompson5910):

the 1080/1296 represents the cases when a does not equal c the 1/6 represents when a = c (d = b would also have to be true)

OpenStudy (rizags):

wait, why 5 times 6^3 why 5?

jimthompson5910 (jim_thompson5910):

because if a doesn't equal c, then either a or c has one less choice. Let's say a = 2 if a doesn't equal c, then c cannot choose 2 to land on. It can only choose from {1, 3, 4, 5, 6} so 5 choices

jimthompson5910 (jim_thompson5910):

normally it is 6^4 or 6*6*6*6 but one of the die has been restricted to 5 choices, so 5*6*6*6 = 5*6^3 = 1080

OpenStudy (rizags):

isnt that dependent on b=d?

jimthompson5910 (jim_thompson5910):

no if a doesn't equal c, then b = d is not important b could equal d, or it could not, we don't care really

jimthompson5910 (jim_thompson5910):

b = d is only crucial when a = c

jimthompson5910 (jim_thompson5910):

because if a = c and \(b \ne d\) leads to a problem of division by zero leading to no solutions at all

OpenStudy (rizags):

what are the chance of b=d AND a=c occuring

jimthompson5910 (jim_thompson5910):

1/6 times 1/6

OpenStudy (rizags):

so the answer can be defined as \[\frac{ 5\times6^{3} }{ 6^{4} }+\frac{ 1 }{ 36 }\]

OpenStudy (rizags):

Why does that work?

OpenStudy (rizags):

Ahhh, is it because we are adding the total amount of ways that a is not equal to c, with the total amount of ways that a=c AND b=d?

OpenStudy (rizags):

does that sum up every possibility?

jimthompson5910 (jim_thompson5910):

they made a typo: on the solution 1, it *should* say "adding 1/36 to 1080/1296" so you are correct because: 1080/1296+1/36 = 31/36 while: 1080/1296+1/6 = 1

jimthompson5910 (jim_thompson5910):

but yes you have the correct idea

OpenStudy (rizags):

ok, since i need to use this for a brief paper, i need to generalize it.

OpenStudy (rizags):

would it be fair to say, that given two parabolas with equations y= x^2 + ax +b and y= x^2 + cx +d, where a, b, c, and d are integers, each chosen independently by rolling a fair n-sided die, the probability that the graphs of the two intersect is equal to \[1-\frac{ 1 }{ n}\times \frac{ n-1 }{ n } = \frac{ n^{2}-n+1 }{ n^{2} }\]

jimthompson5910 (jim_thompson5910):

yes you can use solution 2 to generalize as it might be easier using solution 1 gives you \[\Large \frac{(n-1)*n^3}{n^4} + \frac{1}{n^2}\] \[\Large \frac{(n-1)*n^3}{n^4} + \frac{n^2}{n^4}\] \[\Large \frac{(n-1)*n^3+n^2}{n^4}\] \[\Large \frac{n^4-n^3+n^2}{n^4}\] \[\Large \frac{n^2(n^2-n+1)}{n^4}\] \[\Large \frac{n^2-n+1}{n^2}\] so either way works

OpenStudy (rizags):

ok great, thanks. One more thing though, would it be extremely tedious (and maybe impossible) to generalize this for an n-th degree equation?

OpenStudy (rizags):

that was my ultimate goal, but I'm not so sure if its possible any more

jimthompson5910 (jim_thompson5910):

so instead of quadratics, any two polynomials of degree n?

OpenStudy (rizags):

right

jimthompson5910 (jim_thompson5910):

well the nice thing about the quadratics one we were given is that the quadratic portion cancelled out leaving us with linear equations. Quadratics are hard enough, cubics are even more impossible (still doable). Same can be said about quartics. according to this, quintics can't even be solved exactly http://mathworld.wolfram.com/QuinticEquation.html and the same is true for any degree larger than 5

jimthompson5910 (jim_thompson5910):

sure you can get numerical approximations, so that's at least something. But basically I'm trying to say that it gets even more complicated as n increases

OpenStudy (rizags):

well, would it be fair to say in my paper, that for any n-th degree polynomial that maintains just the four variables at a, b, c, and d, Such as the equations y=x^5 + x^4 + x^3 + x^2 + ax+ b y= x^5 + x^4 + x^3 + x^2 + cx+ d the proability will be (n^2 - n +1)/n^2, dependent on the n-value of the die?

jimthompson5910 (jim_thompson5910):

oh you're still sticking to that format, I see

OpenStudy (rizags):

because no matter how high the degree of the polynomial, the x^m powers can be subtracted off to get the original ax +b = cx+d?

OpenStudy (rizags):

is that true?

jimthompson5910 (jim_thompson5910):

so all that other extra complicated stuff would cancel

jimthompson5910 (jim_thompson5910):

yes that looks good

OpenStudy (rizags):

i suppose that the paper would be better if i tried an equation where one of the variable was behind an x^2, but i dont think i could solve it, so i will stick to this simplified generalization

jimthompson5910 (jim_thompson5910):

so yeah, you can say "the \(\large x^m\) terms cancel where \(\large 2\le m \le n\)"

jimthompson5910 (jim_thompson5910):

yeah I don't know how you'd solve it that way

OpenStudy (rizags):

wait, why must you include the n in the inequality?

OpenStudy (rizags):

i could have an 8th degree polynomial with a 6 sided die couldnt i?

jimthompson5910 (jim_thompson5910):

I put in the n to restrict how high m goes

jimthompson5910 (jim_thompson5910):

when you are referring to the nth degree polynomials

OpenStudy (rizags):

but why does it need to be restricted

jimthompson5910 (jim_thompson5910):

because if n = 12, then m = 13 isn't possible

jimthompson5910 (jim_thompson5910):

if n = 12, all the terms from x^2 to x^12 will cancel

OpenStudy (rizags):

wait, n is not referring to the sides of the die is it?

jimthompson5910 (jim_thompson5910):

no, the two sides of the equation

jimthompson5910 (jim_thompson5910):

the degrees of the two polynomials

jimthompson5910 (jim_thompson5910):

oh wait, you used n to be the sides of the die

OpenStudy (rizags):

ok, i understand, i thought you were referring to n- sides of the die, as in the previous generalization. I understand now

jimthompson5910 (jim_thompson5910):

hmm, then you'll have to use another variable for the generalized polynomials

OpenStudy (rizags):

so for a z-th degree polynomial, what would my inequailty be?

jimthompson5910 (jim_thompson5910):

\[\Large 2 \le m \le z\]

OpenStudy (rizags):

ok, so the x^m terms cancel, when m is greater than or equal to 2, and less than or equal to the degree of the polynomial?

jimthompson5910 (jim_thompson5910):

you'd say "the \(\large x^m\) terms cancel where \(\large 2\le m \le z\)" or something like that

OpenStudy (rizags):

got it, great Thanks so much for your help. By the way, do you know what a mathematical paper abstract should look like given this: "The abstract is fundamental. First, it identifies the subject and it gives details that didn't fit into the title. Then, it describes the main issues and briefly mentions the discussions included in the paper. It should not include any general or background information. It can be viewed as a table of contents, but written in prose."

OpenStudy (rizags):

like for this problem what should i put in the abstarct?

jimthompson5910 (jim_thompson5910):

The abstract is basically what you're trying to do in the entire paper. It's like a summary of the entire paper for those who really don't want to read all of it. It's helpful to save time when you're conducting research. If you're trying to learn more about topic X, and topic X is mentioned in the abstract, then you're probably going to read the paper. Ideally you'd find your answer, but that doesn't happen every time. If you don't get your answer, at least you might learn some cool new trick or property or something.

OpenStudy (rizags):

so basically, it needs to be brief and to the point?

jimthompson5910 (jim_thompson5910):

yeah relatively short

jimthompson5910 (jim_thompson5910):

So in your case, you're going to describe the problem. In addition, you're going to briefly mention how you found the solutions for the particular and general cases. Of course, the actual full descriptions will be in the body of the paper and not in the abstract.

OpenStudy (rizags):

so no lengthy mathematical equations in the abstract?

jimthompson5910 (jim_thompson5910):

No, just the basic idea of them. The abstract is about a paragraph in length I think. Usually it's all words.

OpenStudy (rizags):

ok, got it thank you very much i appreciate your help

jimthompson5910 (jim_thompson5910):

you're welcome